40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jul 19 18:41:42 2022, yuta, Update, BHD, Contrast measurements for Michelson and ITM-LO ContrastMeasurements.pdf
    Reply  Wed Jul 20 18:07:52 2022, Paco, Update, BHD, BHD MICH test BHD_MICH_OSC.pdf
       Reply  Tue Jul 26 20:54:08 2022, Paco, Update, BHD, BHD MICH test - LO phase control 20220726_BICHD.pdf
          Reply  Wed Jul 27 18:30:50 2022, yuta, Update, BHD, LO beam power at BHD DCPDs is significantly lower than expected LOBeamAtBHD.JPGITMXSingleBounceAtBHD.jpg
             Reply  Thu Jul 28 14:34:40 2022, Yehonathan, Update, BHD, LO beam power at BHD DCPDs is significantly lower than expected 
                Reply  Thu Jul 28 16:51:55 2022, Tega, Update, BHD, Shaking test for LO beam AS beam to BHD DCPDs Screenshot_2022-07-28_16-58-34_LOandASShaking.pngScreenshot_2022-07-28_18-08-55_DCPDPOPSuspensionCoherence.png
                   Reply  Thu Jul 28 20:16:26 2022, Anchal, Update, BHD, Shaking test for LO beam AS beam to BHD DCPDs 
                   Reply  Fri Jul 29 18:24:53 2022, Tega, Update, BHD, LO beam power improved by factor of 6 after LO and AS beam alignment LO_and_oldAS_settings.pngBHD_fringe_settings.pngScreenshot_2022-07-29_19-37-39.pngFromTheLeft-AS-POP-LO.JPG
                      Reply  Wed Aug 3 16:00:51 2022, yuta, Update, BHD, BHD fringe aligned with reduced LO and AS beam clipping Screenshot_2022-08-03_15-46-36_BHDfringeAlmostUnclipped.png
                         Reply  Tue Aug 9 15:33:12 2022, yuta, Update, BHD, BHD fringe contrast improved from 43% to 74% HPC-DCPD_B_OUT_1344118517_ITMY-LO.pngHPC-DCPD_A_OUT_1344118517_ITMY-LO.pngHPC-DCPD_B_OUT_1344118318_ITMX-LO.pngHPC-DCPD_A_OUT_1344118318_ITMX-LO.png
                            Reply  Tue Aug 9 15:50:22 2022, Koji, Update, BHD, BHD fringe contrast improved from 43% to 74% 
                               Reply  Wed Nov 16 15:09:08 2022, yuta, Update, BHD, BHD fringe contrast measured with unwhitening filters ContrastMeasurements20221116_edited.pdf
                         Reply  Tue Oct 25 08:25:00 2022, JC, Update, BHD, BHD fringe aligned with reduced LO and AS beam clipping 
                         Reply  Thu Nov 17 11:24:39 2022, JC, HowTo, LSC, Locking MICH 
                            Reply  Tue Nov 29 11:38:37 2022, JC, HowTo, LSC, Lock Single Arm After MICH lock Screen_Shot_2022-11-29_at_11.01.56_AM.png
Message ID: 17020     Entry time: Tue Jul 19 18:41:42 2022     Reply to this: 17024
Author: yuta 
Type: Update 
Category: BHD 
Subject: Contrast measurements for Michelson and ITM-LO 

[Paco, Yuta]

We measured contrast of Michelson fringe in both arms locked and mis-aligned. It was around 90%.
We also measured the contrast of ITM single bounce vs LO beam using BHD DC PDs. It was around 43%.
The measurement depends on the alignment and how to measure the maximum and minimum of the fringe. ITM-LO fringe was also not stable because motions of AS/LO mirrors are large. More tuning necessary.

Background
 - As measured in elog 40m/17012, we see a lot of CARM in AS, which indicates large contrast defect.
 - We want to check mode-matching of LO beam to AS beam.

BHD DC PD conditioning
 - We added DCPD_A and DCPD_B to /opt/rtcds/caltech/c1/scripts/LSC/LSCoffsets3 script, which zeros the offsets when shutters are closed.
 - We also set C1:LSC-DCPD_(A|B)_GAIN = -1 since they are inverted.

Contrast measurement
 - Contrast was measured using channels ['C1:LSC-ASDC_OUT','C1:LSC-POPDC_OUT','C1:LSC-REFLDC_OUT','C1:LSC-DCPD_A_OUT','C1:LSC-DCPD_B_OUT']. For LO, only DCPD_(A|B) are used.
 - We took 15%-percentile (40% for ITM-LO fringe) from the maximum and minimum of the data, and took the median to estimate the maximum value and the minimum value (see Attachment).
 - Contrast = (Imax - Imin) / (Imax + Imin)
 - We measured three times in each configuration to estimate the standard error.
 - Jupyter notebook: https://git.ligo.org/40m/scripts/-/blob/main/CAL/BHD/measureContrast.ipynb

Results
Both arms locked, MICH fringe (15% percentile)
Contrast measured by C1:LSC-ASDC_OUT is 89.75 +/- 0.17 %
Contrast measured by C1:LSC-POPDC_OUT is 79.41 +/- 0.86 %
Contrast measured by C1:LSC-REFLDC_OUT is 97.34 +/- 0.34 %
Contrast measured by C1:LSC-DCPD_A_OUT is 95.41 +/- 1.55 %
Contrast measured by C1:LSC-DCPD_B_OUT is 89.76 +/- 1.49 %
Contrast measured by all is 90.34 +/- 1.68 %


Both arms mis-aligned, MICH fringe (15% percentile)
Contrast measured by C1:LSC-ASDC_OUT is 89.32 +/- 0.57 %
Contrast measured by C1:LSC-POPDC_OUT is 94.55 +/- 0.62 %
Contrast measured by C1:LSC-REFLDC_OUT is 97.95 +/- 1.37 %
Contrast measured by C1:LSC-DCPD_A_OUT is 96.40 +/- 1.04 %
Contrast measured by C1:LSC-DCPD_B_OUT is 90.98 +/- 1.07 %
Contrast measured by all is 93.84 +/- 0.94 %


ITMY-LO fringe (40% percentile)
Contrast measured by C1:LSC-DCPD_A_OUT is 45.51 +/- 0.45 %
Contrast measured by C1:LSC-DCPD_B_OUT is 38.69 +/- 0.43 %
Contrast measured by all is 42.10 +/- 1.03 %


ITMX-LO fringe (40% percentile)
Contrast measured by C1:LSC-DCPD_A_OUT is 46.65 +/- 0.65 %
Contrast measured by C1:LSC-DCPD_B_OUT is 39.82 +/- 0.51 %
Contrast measured by all is 43.24 +/- 1.45 %


Discussion
 - As you can see from the attachment, REFLDC is noisy and over estimating the contrast. ASDC is reliable. We need to tune the threshold to measure the maximum value and minimum value. We should also use the mode instead of median.
 - Contrast depends very much on the alignment. We didn't tweak too much today.
 - ITM-LO fringe was not stable, probably due to too much motion in AS1, AS4, LO1, LO2. Their damping needs to be re-tuned.

Next:
 - Model FPMI sensing matrix with measured contrast defect
 - Estimate AS-LO mode-mismatch using the measured contrast
 - Lock ITM-LO fringe using DCPD_(A|B) as error signal, and ITM or LO1/2 as actuator
 - Lock MICH with DCPD_(A|B), and with LO beam
 - Get better contrast data with better alignment and better AS1, AS4, LO1, LO2 damping

Attachment 1: ContrastMeasurements.pdf  178 kB  Uploaded Tue Jul 19 20:29:13 2022  | Hide | Hide all
ContrastMeasurements.pdf ContrastMeasurements.pdf ContrastMeasurements.pdf ContrastMeasurements.pdf ContrastMeasurements.pdf ContrastMeasurements.pdf ContrastMeasurements.pdf ContrastMeasurements.pdf
ELOG V3.1.3-